# Explainable Deep Learning Analysis for Raga Identification in Indian Art Music

### v1

@@ V-B4 CN2+LSTM+T
Figure 4: Raga class distribution for test files
V-C Saliency Evaluation
V-C1 Saliency annotation
@@
The music expert is given a randomly selected 64 audio files for which the model predicts the correct class. The task is to look for such regions in the audio file which would allow any hindustani music expert to judge if that audio snippet belongs to a particular Raga class. The Expert is asked to keep only those audio files which have at least 5 to 7 seconds of such region and at max 15 to 17 seconds. As it would make no sense to mark just one or two seconds, or the whole audio as the salient region. The salient region can be a single region, or may be 2 or 3 disjoint regions throught the 30 second audio chunk. Also, it was made sure that we have at least 3 audio examples per class. So after the annotation, we get 47 labelled test examples to evaluate our saliency methods. The distribution of test classes is shown in figure 4.
V-C2 Saliency evaluation
To evaluate the salient regions identified by the two ExAI models, we compare the predictions with the human-annotated saliency regions. For GC, we sum the activation map values across the frequency to get a 938x1 vector, showing saliency values for all 938 frames. We also convert the annotations of music experts into a multi-hot vector of length 938, where we mark 1 for salient region and 0 for non-salient. Now, to compare, we try out different windows starting from top 1 seconds upto top 10 seconds of GC prediction values and compute precision in each case.
For SL model, we get coefficients of linear regression model trained on the test data, in terms of 30 masks. We upscale this to get a multi-hot vector of 938x1 to match the dimensions of expert annotations. Now we compare the salient region given by this model in the same way as GC and compute precision values.
In both these cases, we compute the model’s precision for upto top 10 seconds out of the salient region, starting from 1 second. This is done so that we are able to judge how correct the top salient region of our model is. If the model is giving out decent results, it can be used as snippet for being representative of the particular Raga to use that for music teaching other applications.
TABLE III: Saliency report for GC and SL
@@ V-C2 Saliency evaluation
V-C3 Saliency results
Table: III shows the achieved average precision of explanations over the 47 test samples, compared with expert annotations for the most-salient N-seconds for both the ExAI models, where $N\in[1:10]$. Observe that for a smaller N, precision is high and it decreases as N increases, indicating that the top N seconds of highlighted salient regions by these ExAI models are very close to actual salient regions for that particular audio chunk, implying that the model is indeed relying on the time frames (the music in those time-frames) to make its predictions for current Raga class. Also, we can observe that SL shows higher performance which is also expected since GC considers activations only upto the CNN layers, leaving out the LSTM layers. Although there is no direct comparison metric among the two methods, we argue here that SoundLIME explanations will be more reliable as it takes predictions from the whole model (including the LSTM layers) while generating the explanations, hence these should prove to be a better representative of the predictions of the classification model.
TABLE IV: Analysis of individual samples
@@ V-C3 Saliency results
Now, we study two individual test audios for salient regions. We select a ramdom correctly predicted test sample belonging to Raga class “Des”. The classifier shows a softmax probability score of 0.95 as shown in Table: IV with signifying that the model is quite confident about this prediction. The true salient region marked by the music expert, the salient region as reported by GC and SL are shown in Figures 5. The marked true region by the music expert came out to be around 15 seconds. We are concerned about finding how much of the region predicted by the model falls in the correct category. The precision score for GC and SL come out to be 0.77 and 0.8 respectively for 10 seconds, signifying a significant overlap and hence a significant amount of the prediction being based on the true concepts related to music only.
@@
We also conduct an individual study on two closely related Ragas in HCM, namely Bihag and Maru-Bihag. Maru-Bihag, being a musical blend of two Ragas, ”Maru” and ”Bihag”, may exhibit some characteristics of Raga Bihag in its composition, and vice versa. We can observe in the confusion matrix shown in Fig: 7 also that there are a few number of samples of class Bihag which have been mis-classified as Maru-Bihag.
Figure 7: Confusion Matrix for 12-class classifier
We carefully select such a false negative example and get it annotated for salient region showing the current Raga (Bihag in this case). Fig: 6 shows the true salient region, predicted salient region from GC and SL. We obtained top k seconds of salient region to match the length of salient region obtained from expert annotations from GC and SL models, shown in Fig: 6. The softmax probabilities as shown in table:IV of this model are 0.26 (true class: Bihag) and 0.71 (predicted class: Maru-Bihag), signifying that the model does recognise the true class (Bihag) to some extent in the given audio, but prediction of Maru-Bihag suggests that there is some ambiguous region in that particular audio snippet which confuses the model to differentiate between the two Raga classes. The same is also evident from the fact that the expert was only able to find a very small region from the whole audio that prominently shows the characteristics of Raga Bihag, implying that the audio chunk itself could be a bit ambiguous. Also, The precision score for GC and SL come out to be 0.12 and 0.0 respectively. We can clearly see in the figure6 that the salient regions of both GC and SL do not overlap the true region and points to totally different region, implying that there were frames which are similar to Raga Maru-Bihag in that audio sample and the frames implying Raga Bihag might not be that prominent, which can be confirmed by listening to the audio also.
@@ VI Conclusion and Future Scope
In this paper, we introduce a novel labeled dataset for Hindustani Classical Music (HCM) analysis and achieve an f1-score of 0.89 for 12-class Raga classification task by carrying out Ablation studies. We observe that, providing tonic normalization to the extracted Chromagram features significantly enhances the performance of the Raga classifier. Following that, we employ two Explainable AI (ExAI) models to identify the salient regions for the classifier’s predictions and analyze their overlap with human annotations. This comparison reveals a significant overlap, affirming that the patterns or characteristics theoretically associated with the identification of a Raga for any audio are present and are being used by the Raga Identification model for the task. This serves as a correctness check for the music theory that specific motifs, phrases or characteristics are indeed responsible for the identification of any Raga by the Deep Learning model built on a dataset that represents actual HCM performances.
This work provides a valuable contribution to the field of explainability in Raga identification in Indian Art Music, setting the foundation for future research. The meticulously labeled dataset, including Ragas and tonic information, expands the landscape of available high-quality datasets in HCM analysis, and opens avenues for further research in music classification and analysis. Moving forward, Incorporating musicians’ perspectives and musical concepts like ”Khatka” and ”murki” to refine Raga classifiers can ensure correctness of Raga identification according to musical conventions and nuances.
Future research can focus on developing methods to generate listenable or musical explanations, which could serve as standalone identifiers of the Raga. This would include extracting phrases from audio clips that correspond to the Raga, offering potential applications in music pedagogy and beyond. It is important to note that experts may not always agree on the same set of phrases or regions, indicating the subjective nature of some aspects of music analysis. Another direction for future research can be to extend this work to other musical traditions and explore the generalizability of the approach.

### v2

@@ V-D Train on the whole dataset
Figure 4: Visualisation of Raga class distribution for test files selected for XAI analysis
V-E Saliency Evaluation
Now, we conduct an in-depth analysis of the salient regions identified by the XAI models to understand their musical significance and validate the manual annotations provided by the music expert.
V-E1 Saliency annotation
@@
The music expert is given a randomly selected set of 64 audio chunks out of $S_{t}$ for which the model predicts the correct class. The task is to look for regions in each audio file that contain musical elements such as the Pakad, Aaroh, Avroh, or Mukh-Ang of the raga: features that would allow any HCM expert to determine if the audio snippet belongs to a particular raga class. The expert selects only those audio files that have at least 5 to 7 seconds, and at most 15 to 17 seconds of such salient regions as it makes little sense to mark just one or two seconds, or the entire audio, as the salient region. The salient region can be a single segment or may consist of two or three disjoint regions throughout the 30-second audio chunk. We also ensure that we have at least three audio examples per class for all $N=12$ classes. After the annotation process, we obtain 47 labeled test examples $S_{ts}\subset S_{t}$ to evaluate our saliency methods. The distribution of test classes is shown in Fig. 4.
V-E2 Saliency evaluation
We compare the salient regions identified by GC and SL with expert annotations to evaluate the models’ ability to highlight musically significant regions. For GC, we sum the activation map values across the frequency axis to obtain a saliency vector over time frames. This vector indicates the importance of each time frame in the model’s prediction. For SL, we obtain the coefficients from the linear regression model, representing the importance of each of the 30 time masks. We upsample these coefficients to match the time resolution of the expert annotations. We calculate the precision of the models’ salient regions against the expert annotations for the top N seconds (N = 1 to 10), assessing how well the most critical regions identified by the models align with musically significant parts of the audio.
TABLE IV: Saliency report for GC and SL
@@ V-E2 Saliency evaluation
Figure 5: Precision vs. Accuracy: The graph shows the relationship between the precision of the XAI model (binned with increments of 0.5) and the classification model’s accuracy. The yellow circles represent the weight of each bin, indicating the number of points in each bin.
V-E3 Saliency results
Table IV presents the average precision for both the XAI models over the test-set $S_{ts}$, compared with expert annotations. The precision values are calculated for the top $t$-second ($t\in[1:10]$) regions, where $t=1$ represents how well the most salient 1-second region predicted by the XAI model aligns with the salient region annotated by the expert. Observe that for a smaller t, precision is high and it decreases as t increases, indicating that the top t seconds of the salient regions highlighted by these XAI models are very close to actual salient regions for that particular audio chunk, implying that the model is indeed relying on the appropriate time frames (the music in those time-frames) to make its predictions for current Raga class.
Let us now observe the relationship between the precision of the XAI model and the accuracy of the classification model. We bin the precision scores of the XAI model with an increment of 0.5. For each bin, we compute the classification model’s accuracy and plot Precision vs Accuracy as shown in Fig. 5. We observe a clear trend showing that higher precision from the XAI model correlates with higher classification accuracy. This indicates that when the XAI model highlights regions with high precision, the classifier is more likely to make correct predictions, indicating that the generated explanations are indeed meaningful and align with human understanding.
From Table IV, we observe that SL outperforms GC, particularly for the top few seconds. This behavior is expected, as GC considers activations only from the CNN layers, excluding the contributions of the LSTM layers. While there is no direct metric to compare the two methods, we argue that SL provides more reliable explanations, as it incorporates predictions from the entire model, including the LSTM layers, making it a better representation of the classification model’s decision-making process in our case.
TABLE V: Analysis of individual samples
@@ V-E3 Saliency results
V-F Musical Analysis of Salient Regions
V-F1 Example 1: Raga Des
For comparison, we analyze a randomly selected, correctly predicted sample of Raga Des, where the classifier model confidently assigns a softmax probability of 0.95 as shown in Table V. For the audio, the expert was able to identify a salient region of about 16 seconds, which includes the note sequence: ‘Pa Dha me Ga Re ma Ga Re Ga Ni Sa’, which resembles the pakad (signature phrase) of Raga Des. Both GC and SL highlight regions that overlap significantly with this segment annotated by the expert, with precision scores of 0.77 and 0.80 respectively for the top 10 seconds, as shown in Table V. This significant overlap suggests that the XAI models are looking for musically meaningful regions to make their predictions for salient regions, and are aligning well with human understanding of Raga as a concept.
@@ V-F1 Example 1: Raga Des
V-F2 Example 2: Raga Bihag misclassification
We now conduct a study on two closely related Ragas in HCM, namely Bihag and Maru-Bihag. Maru-Bihag, being a musical blend of two Ragas, “Maru” and “Bihag”, is supposed to exhibit some characteristics of Raga Bihag in its composition. We can observe in the confusion matrix shown in Fig. 8 also that there are a few samples of class Bihag that have been misclassified as Maru-Bihag.
Figure 8: Confusion Matrix for predictions on test dataset for CN2+LSTM+T model
We analyze a false negative example where the model incorrectly classifies the true Raga Bihag as Maru-Bihag. The annotated salient region for Bihag, along with the predicted regions from GC and SL, is shown in Figure 7. The softmax probabilities (Table V) indicate 0.26 for the true class (Bihag) and 0.71 for the predicted class (Maru-Bihag), suggesting that the model partially recognizes the true class but is misled by overlapping characteristics between the two Ragas present in the test audio.
Expert annotations reveal that only a small portion of the audio distinctly represents Raga Bihag (Fig. 7), indicating inherent ambiguity in the audio itself. This is further reflected in the low precision scores for GC (0.12) and SL (0.0) in Table V, as their predicted salient regions fail to align with the annotated region and instead highlight unrelated sections of the audio. The mismatch suggests that the audio contains regions with characteristics closer to Maru-Bihag, while the distinctive features of Bihag are less prominent. This observation underscores the complexity of distinguishing between closely related Ragas and is corroborated by expert listening and annotation.
@@ VI Conclusion and Future Scope
In this paper, we curate a dataset for Hindustani Classical Music analysis from Prasar-Bharti audios, manually annotate it with Raga and Tonic labels, and achieve an f1-score of 0.89 for the 12-class Raga classification task using the proposed model. We observe that providing tonic normalization to the chromagram features significantly enhances the performance of the Raga classifier. Following that, we employ two Explainable AI (XAI) models: GradCAM++ and SoundLIME, to identify the salient regions for the classifier’s predictions and analyze their overlap with human annotations. Our comparison reveals a significant overlap, affirming that the patterns or characteristics theoretically associated with the identification of a Raga for any audio are present and are being used by the Automatic Raga Identification (ARI) model for the task. This serves as a correctness check for the music theory that specific motifs, phrases or characteristics are indeed responsible for the identification of any Raga by the Deep Learning model built on a dataset that represents actual HCM performances. It is important to note that experts may not always agree on the same set of phrases or regions, indicating the subjective nature of such music analysis.
This work provides a valuable contribution to the field of explainability in ARI for Indian Art Music, setting the foundation for future research. Moving forward, incorporating musicians’ perspectives and musical concepts like “Khatka” and “murki” to refine Raga classifiers can ensure correctness of ARI according to musical conventions and nuances. Future research can focus on developing methods to generate listenable or musical explanations, which could serve as standalone identifiers of Raga. This can include extracting phrases from audio clips that correspond to the Raga, offering potential applications in music pedagogy and beyond.